OpenStudy (anonymous):

find the following limit [(1+2x)^(1/3)-1]/x lim->0

OpenStudy (anonymous):

\[\huge \lim_{x\to 0}\frac{\sqrt[3]{1+2x}-1}{x}\] I rewrote it better.... you're supposed to know this formula: \[\Large a^3-b^3=(a-b)(a^2+ab+b^2)\] ...

OpenStudy (turingtest):

that's not right kreshnick, you want to normalize the numerator

OpenStudy (anonymous):

Thanks for writing it in good form though

OpenStudy (anonymous):

@TuringTest what's wrong O_O :(

OpenStudy (turingtest):

well I don't see how to do it that way, let's see what you were gonna do

OpenStudy (anonymous):

alright... do you have patience to wait till I'll write it with latex?

OpenStudy (turingtest):

sure I gotta bounce for a bit anyway, you guys work it out and I'll be happy to see when I get back

OpenStudy (anonymous):

alright then ... I might be wrong... but I'll show what I meant

OpenStudy (turingtest):

it's even better when I'm wrong, then I get to learn new ways to do things :)

OpenStudy (turingtest):

for the record my thought is to multiply by\[{(1+2x)^{2/3}+1\over(1+2x)^{2/3}+1}\]but whatever works :)

OpenStudy (anonymous):

sorry to interrupt...is my question right now? (TuringTest)

OpenStudy (turingtest):

the way kreshnick has it is the same as yours, if that is what you mean

OpenStudy (turingtest):

oh heck, if you can use l'hospital you can do it in a flash

OpenStudy (turingtest):

ok brb

OpenStudy (anonymous):

without L'hospital please

OpenStudy (anonymous):

\[\Large \lim_{x\to 0}\frac{\sqrt[3]{1+2x}-1}{x}\cdot \frac{\left[(\sqrt[3]{1+2x}-1)^2+(\sqrt[3]{1+2x}-1)+1^2\right]}{\left[(\sqrt[3]{1+2x}-1)^2+(\sqrt[3]{1+2x}-1)+1^2\right]}\] \[\Large \lim_{x\to 0}\frac{(\sqrt[3]{1+2x})^3-1^3}{x\left[(\sqrt[3]{1+2x}-1)^2+(\sqrt[3]{1+2x}-1)+1^2\right]}\] \[\Large \lim_{x\to 0}\frac{1+2x-1}{x\left[(\sqrt[3]{1+2x}-1)^2+(\sqrt[3]{1+2x}-1)+1^2\right]}\] \[\Large \lim_{x\to 0}\frac{2x}{x\left[(\sqrt[3]{1+2x}-1)^2+(\sqrt[3]{1+2x}-1)+1^2\right]}\] \[\Large \lim_{x\to 0}\frac{2}{\left[(\sqrt[3]{1+2x}-1)^2+(\sqrt[3]{1+2x}-1)+1^2\right]}\] \[\Large =\frac{2}{\left[(\sqrt[3]{1+2\cdot 0 }-1)^2+(\sqrt[3]{1+2\cdot 0 }-1)+1^2\right]}=\] \[\Large =\frac{2}{\left[(1-1)^2+(1-1)+1^2\right]}=\] \[\Large =\frac{ 2}{1}=2\] @TuringTest I think I deserve a medal !! LOL

OpenStudy (anonymous):

No L'hopital's Rule.. that would be a piece of cake !

OpenStudy (anonymous):

thanks man. I knew the formula but, I didn't know how to apply it in this question.

OpenStudy (anonymous):

Glad to help ... but hold on a second ! @TuringTest still there?

OpenStudy (anonymous):

he is going to come to look at it.

OpenStudy (turingtest):

I'm sorry to say that your answer's not quite right :(

OpenStudy (turingtest):

try it with l'Hospital and you will see

OpenStudy (anonymous):

... ahh , what's wrong :( [ don't be sorry, I'm glad to know that I can learn something ]

OpenStudy (turingtest):

not sure, I just got back... I have to look it over

OpenStudy (anonymous):

with L'Hopital's Rule I get 2/3 ... O_O

OpenStudy (anonymous):

what do you get ?

OpenStudy (turingtest):

the answer is 2/3

OpenStudy (turingtest):

I see that you were inconsistent in your use of a in your formula you made a=(1+2x)^(1/3) at times, and a=(1+2x)^(1/3)-1 at others I see the logic in your work though, you got the top number right :D

OpenStudy (turingtest):

arg, I can't retype the difference in your latex line....

OpenStudy (anonymous):

I used .. \[\LARGE a-b= \sqrt[3]{1+2x}\] so I had to multiply by \[\LARGE a^2+ab+b^2\] to get the third root .....

OpenStudy (anonymous):

I GET IT..... I put there -1 too ... pfff... but if that wouldn't be there THIS would be CORRECT ... in the end we get.. \[\LARGE \frac{2}{1+1+1}\]

OpenStudy (turingtest):

I though you were saying \(a=\sqrt[3]\)\[\lim_{x\to 0}\frac{\sqrt[3]{1+2x}-1}{x}\cdot \frac{\left[(\sqrt[3]{1+2x})^2+\sqrt[3]{1+2x}+1^2\right]}{\left[(\sqrt[3]{1+2x}^2+(\sqrt[3]{1+2x})+1^2\right]}\]remember that a= ...and I can see you figure out your mistake nice technique though, new one on me :)

OpenStudy (anonymous):

so I just made a typing mistake, this way of doing it works too !

OpenStudy (turingtest):

I like being wrong :)

OpenStudy (anonymous):

at least I knew what I was doing LOL hahahaha... (ME TOO !)